OpenStudy (anonymous):

5.2 = 2(0.7)^2 + 4(0.7) h - how can i get the "h" in that equation. that was 2(0.7) squared. just dunno how to do the squared thing here :) please help. ;)

OpenStudy (anonymous):

This is what you wanted to write. The equation editor is your friend. :) \[5.2 = 2(0.7)^{2} + 4(0.7)h\] First combine your constants that are in the same terms. \[2(0.7)^{2} = 2 * 0.7 * 0.7 = 0.98\] \[4(0.7) = 2.8\] And rewrite the equation as \[5.2 = 0.98 + 2.8h\] Can you go on from there?

OpenStudy (anonymous):

is the answer h= 1.50? is it right? :/

OpenStudy (anonymous):

It's very close to 1.5 but not exactly, because 4.22 / 2.8 does not come out even. Do you need an exact answer, or just an approximate number?

OpenStudy (anonymous):

the exact answer please. so 1.5 could be the answer already?

OpenStudy (anonymous):

No. If you need an exact answer, then you either need to write it as a fraction \[4.22/2.8\] or write the complete decimal, which goes on forever: 1.507142857142857142857142857142857142857142857142857142857142857142... Writing forever will take a very long time though. :)

OpenStudy (anonymous):

oh yea. so i mean i need the estimated one. 1.50 would do :) thanks! :)

OpenStudy (anonymous):

lol. I was hoping you would not write the whole decimal. Though you could make a youtube video of yourself doing that and maybe become the next internet sensation.

OpenStudy (anonymous):

hahha. yea thanks !! :)